Solve the given integral equation or integro-differential equation for y(t). t y(t) + 27 [(t-vjy(v) dv=t, y(0) =0 0 y(t) =
Answer #1

* {s = $33} - ੩ (6) $ 5 * 11:31. 341 1 + 8 = f ACE - Z) gcoida = - 1 (E ) + 27 fce v) y Nav t (c ) 0 f(t) = tu 9 CE) : ੧ ) S